please solve, show work and simplify
1. t(x)=3x-(x2-1)
g(x) =2x+3
Find: (tog)(x) is it function or not
THANKS!!!!!
redrobyn47

Answer by Nate(3500)

(t o g)(x) means t[g(x)]
t(x) = 3x - x^2 + 1
t(2x + 3) = 3(2x + 3) - (2x + 3)^2 + 1
t(2x + 3) = 6x + 9 - 4x^2 - 12x - 9 + 1
t(g(x)) = -4x^2 - 6x + 1